# Topology optimization of nonlinear forced response curves via reduction on spectral submanifolds

### v1

@@ 2.3.2 Control of SN bifurcations on FRC
As discussed in li_nonlinear_2022 ; liang_bifurcation_2025 , the SN bifurcations of the fixed points of the reduced dynamics (15) correspond to those of the periodic orbits of the full system (1). This correspondence allows us to control the occurrence of SN bifurcations in reduced coordinates. In this subsection, we derive the expression of the SN bifurcation points in the reduced coordinates and eliminate their occurrence by enforcing the degeneracy condition of a cusp bifurcation. In the final step, we formulate an optimization problem to suppress the occurrence of SN bifurcations.
We first need to find the fixed point of SN bifurcation $(\rho_{\mathrm{SN}},\Omega_{\mathrm{SN}})$ in reduced coordinates. The Jacobian of the reduced dynamics (15) can be expressed as li_explicit_2024
@@
At the SN bifurcation of a fixed point, the Jacobian $\boldsymbol{A}$ has a simple zero eigenvalue kuznetsov_elements_2023 . Thus, the determinant of $\boldsymbol{A}$ is 0 at SN bifurcation, namely
@@
Note that there are usually two SN bifurcation points on the FRC. Next, we aim to suppress the occurrence of these two SN bifurcation points by enforcing the cusp bifurcation degeneration condition, namely kuznetsov_elements_2023
$b=\langle\boldsymbol{\psi}_{\mathrm{SN}},\ \boldsymbol{B}_{\mathrm{SN}}(\boldsymbol{\phi}_{\mathrm{SN}},\boldsymbol{\phi}_{\mathrm{SN}})\rangle=0.$ (28)
where $\boldsymbol{\psi}_{\mathrm{SN}}$ and $\boldsymbol{\phi}_{\mathrm{SN}}$ are the left and right eigenvectors of $\boldsymbol{A}_{\mathrm{SN}}=\boldsymbol{A}(\rho_{\mathrm{SN}},\Omega_{\mathrm{SN}})$, $\boldsymbol{B}_{\mathrm{SN}}\in\mathbb{R}^{2\times 2\times 2}$ is the second derivative tensor of reduced dynamics (15) at $(\rho_{\mathrm{SN}},\Omega_{\mathrm{SN}})$. Specifically, we have
$\boldsymbol{A}_{\mathrm{SN}}\boldsymbol{\phi}_{\mathrm{SN}}=0,\quad\boldsymbol{A}^{\mathrm{T}}_{\mathrm{SN}}\boldsymbol{\psi}_{\mathrm{SN}}=0,\quad\boldsymbol{B}_{\mathrm{SN}}(\boldsymbol{\phi}_{\mathrm{SN}},\boldsymbol{\phi}_{\mathrm{SN}})=\begin{pmatrix}\boldsymbol{\phi}_{\mathrm{SN}}^{\mathrm{T}}\boldsymbol{B}_{1}\boldsymbol{\phi}_{\mathrm{SN}}\\ \boldsymbol{\phi}_{\mathrm{SN}}^{\mathrm{T}}\boldsymbol{B}_{2}\boldsymbol{\phi}_{\mathrm{SN}}\end{pmatrix}$, (29)
where
$\boldsymbol{B}_{1}=\begin{pmatrix}6\mathrm{Re}(\gamma)\rho_{\mathrm{SN}}&0\\ 0&\mathrm{Re}(\lambda)\rho_{\mathrm{SN}}+\mathrm{Re}(\gamma)\rho_{\mathrm{SN}}^{3}\end{pmatrix},\,\boldsymbol{B}_{2}=\begin{pmatrix}2\mathrm{Im}(\gamma)-2b_{2,1}/\rho_{\mathrm{SN}}^{2}&b_{2,2}\\ b_{2,2}&b_{2,1}\end{pmatrix}$ (30)
with $b_{2,1}=\mathrm{Im}(\lambda)-\Omega_{\mathrm{SN}}+\mathrm{Im}(\gamma)\rho_{\mathrm{SN}}^{2}$ and $b_{2,2}=-(\mathrm{Re}(\lambda)+\mathrm{Re}(\gamma)\rho_{\mathrm{SN}}^{2})/\rho_{\mathrm{SN}}$. Since $\det(\boldsymbol{A})=0$, the eigenvectors $\boldsymbol{\psi}_{\mathrm{SN}}$ and $\boldsymbol{\phi}_{\mathrm{SN}}$ are not uniquely determined. To ensure uniqueness, we impose the following normalization conditions
$\boldsymbol{\psi}_{\mathrm{SN}}^{\mathrm{T}}\boldsymbol{\psi}_{\mathrm{SN}}=1,\quad\boldsymbol{\psi}_{\mathrm{SN}}^{\mathrm{T}}\boldsymbol{\phi}_{\mathrm{SN}}=1.$ (31)
Many MEMS sensors (e.g., accelerometers, gyroscopes) are designed to operate within the linear regime to avoid hysteresis and excess noise caused by nonlinear effects tiwari_using_2019 . To ensure such behavior, we aim to control the system’s bifurcation structure while enhancing its performance. Specifically, we formulate an optimization problem that maximizes the response amplitude $\rho_{\max}$, subject to a constraint on the parameter $b$ to limit the distance between the two SN bifurcation points. The resulting optimization problem is given as
@@
$\quad 0\leq|b|\leq b_{\text{target}}$
@@
where the constant $b_{\text{target}}$ regulates the distance between the two SN bifurcation points. As $b_{\text{target}}$ approaches zero, these two points coalesce, resulting in a cusp bifurcation. In addition, we enforce a frequency constraint on the X mode to ensure connectivity BONACCORSI2025107865 . A two-sided constraint is also imposed on the structural area to improve the stability of the optimization process.
@@ 3 Sensitivity analysis
In the previous section, we formulated several optimization problems. Since the Method of Moving Asymptotes (MMA) is used to solve it, sensitivity analysis with respect to the design variables is essential for updating the design. In this section, we derive the sensitivity expressions for the key quantities $\omega_{j}$, $\lambda$, $\gamma$, $\tilde{f}$, $\rho_{\max}$, and $b$, which are critical for the optimization process.
The sensitivity of the eigenfrequencies $\omega_{j}$ and eigenvalues $\lambda$ has already been derived in Appendix.C of li_explicit_2024 , namely
@@
Following li_explicit_2024 , the variation of $\gamma$ can be obtained by the adjoint method, which can be expressed as
@@
In the above equation, the damping coefficients $\alpha$ and $\beta$ are fixed, while the damping ratio $\xi$ varies throughout the optimization process. This differs from the formulation in li_explicit_2024 , where the damping ratio remains constant. We follow the tensorial approach in pozzi_topology_2025 to ensure efficient implementation.
@@
Similarly, by taking derivative of (28), the sensitivity of $b$ can be expressed as
$\begin{aligned} b^{\prime}=\boldsymbol{\psi}_{\mathrm{SN}}^{\prime}\boldsymbol{B}_{\mathrm{SN}}(\boldsymbol{\phi}_{\mathrm{SN}},\boldsymbol{\phi}_{\mathrm{SN}})+\boldsymbol{\psi}_{\mathrm{SN}}\boldsymbol{B}_{\mathrm{SN}}^{\prime}(\boldsymbol{\phi}_{\mathrm{SN}},\boldsymbol{\phi}_{\mathrm{SN}})+2\boldsymbol{B}_{\mathrm{SN}}(\boldsymbol{\phi}_{\mathrm{SN}},\boldsymbol{\phi}_{\mathrm{SN}}^{\prime})\end{aligned},$ (37)
where the sensitivities of $\boldsymbol{\psi}_{\mathrm{SN}}$, $\boldsymbol{\phi}_{\mathrm{SN}}$, and $\boldsymbol{B}_{\mathrm{SN}}$ are derived in Appendix B.
@@ 4 SIMP Model for Topology Optimization
In this work, the Solid Isotropic Material with Penalization (SIMP) method bendsoe_material_1999 is employed as the foundation for density-based topology optimization. In the SIMP-based framework, the material properties of the structure are interpolated element-wise based on the physical density variables $\hat{\mu}_{e}$. These densities are used to assemble the global matrices and nonlinear force vectors. Specifically, the global mass and stiffness matrices, as well as the quadratic and cubic stiffness tensors, are computed as follows pozzi_topology_2025 :
@@
To ensure mesh-independent and well-posed topology optimization results, the design variable $\mu\in[0,1]$ is first smoothed using a density filter wang_projection_2011 , defined as
@@
Next, a Heaviside-type projection wang_projection_2011 is applied to obtain the projected density $\bar{\mu}_{e}$, which sharpens the transition between solid and void regions:
@@ 5 Numerical examples
In this section, three numerical examples are presented to show the result of minimizing the peak of FRC, controlling hardening/softening behavior, and suppressing the occurrence of SN bifurcations. In all the examples, the 2-dimensional structures are discretized using 4-node, square elements. Following pozzi_topology_2025 , material properties are specified with the density $\rho=2330\times 10^{-6}\ \rm{ng/um^{3}}$, Young’s modulus $E=148\times 10^{9}\ \rm{Pa}$, and Poisson’s ratio $\nu=0.23$. The plane stress approximation is used, with a thickness of 24 $\rm{\mu m}$.
The matrix assembly of the finite element model and the solution of the optimization problems are carried out using the YetAnotherFEcode package Jain2022YetAnotherFEcode . The FRCs of the optimized layouts are computed using SSMTool Jain2023SSMTool , an open-source software package for the computation of invariant manifolds and their reduced dynamics. In the Sec. 2.2 and 2.3.1, we employed a third-order reduced-order model (15) to obtain the expression of FRC in the reduced coordinates (17). To assess the reliability of the optimized result, we further compute the FRC using SSMTool at higher-order truncations for the optimized structure. If the FRC at higher-order truncations agrees well with that at the third-order truncation, it indicates that the predicted FRC has converged and is hence reliable.
@@
We consider a Messerschmitt-Bölkow-Blohm (MBB) beam with a fixed segment located at its center. Due to structural symmetry, the optimization is performed only on the left half of the beam pozzi_topology_2025 , as illustrated in Fig. 1. The half MBB beam measures 800 $\mathrm{\mu m}$ in length and 100 $\mathrm{\mu m}$ in height, including a fixed region with a length of 100 $\mathrm{\mu m}$. The left end of the beam is fully fixed, and the right end is constrained in the $x$-direction. A periodic excitation is applied at the midpoint of the right edge.
@@
With $\gamma_{\mathrm{target}}=5\times 10^{-5}$, the iterative process for the nonlinear optimization formulation (20) is provided in Fig. 2. The objective $\rho_{\max}$ exhibits an initial increase, followed by a rapid decrease until convergence. Meanwhile, the constraint on $\mathrm{Im}(\gamma)$ is violated initially, and the value of $\mathrm{Im}(\gamma)$ gradually converges and ultimately reaches the target $\mathrm{Im}(\gamma)=\gamma_{\mathrm{target}}$. The abrupt changes observed in the objective $\rho_{\max}$ and $\mathrm{Im}(\gamma)$ during the iteration are caused by the scheduled updates of the penalization exponent $p$ and the projection parameter $\beta$ in the SIMP algorithm. The obtained optimal layout is given in the upper panel of Fig. 3. The computational performance of this iterative process is summarized in Table 1.
@@
Since third-order SSM reduction is employed to compute the FRCs of the optimal layouts, it is necessary to assess the convergence of the SSM-based model reduction. The convergence analysis is presented in Appendix C, which shows that the FRC obtained using third-order SSM reduction agrees well with that obtained using seventh-order SSM reduction under the forcing amplitude $f^{\mathrm{ext}}=2\times 10^{10}\ \mathrm{ng\cdot\mu m/ms^{2}}$.
@@
Tailoring the hardening/softening behavior of the backbone curve in an MBB beam has been discussed in pozzi_topology_2025 . In this subsection, we aim to minimize the response amplitude while simultaneously tailoring the softening–hardening behavior, namely, the concurrent design (20). We further compare this concurrent design with the design given by the reference problem (22) that tunes the backbone only. As shown in Fig. 5, the design domain of the beam has a length of 500 $\mathrm{\mu m}$ and a height of 100 $\mathrm{\mu m}$. A fixed region, occupying 20% of the total beam length, is located at the center and serves as a proof mass. The left end of the beam is fully fixed, while the right end is constrained in the $x$-direction. A periodic excitation is applied at the midpoint of the right edge.
@@
The FE model uses a mesh of $100\times 20$ elements, leading to a system with nn= 4179 DOF in (1). The damping ratio of initial layout $\xi$ = 0.1%, which leads to the values of damping constants $\alpha=6.90$ and $\beta=7.41\times 10^{-8}$. During the optimization iteration process, the forcing amplitude is given as $f^{\mathrm{ext}}=5\times 10^{9}\ \mathrm{ng\cdot\mu m/ms^{2}}$. In the following computation, we set $\epsilon=0.01$ unless otherwise stated. The projection parameter is fixed at $\sigma=10$ throughout the optimization, and the filter radius is set to $R=3$. In this example, no constraint is imposed on $\omega_{X}$ since $\omega_{X}>3\omega_{Y}$ is naturally satisfied throughout the optimization iterations. Furthermore, the two-sided constraint on $\gamma$ is replaced by a one-sided constraint. Specifically, $\gamma>\gamma_{\mathrm{target}}$ is enforced when $\gamma_{\mathrm{target}}>0$, and $\gamma<\gamma_{\mathrm{target}}$ is enforced when $\gamma_{\mathrm{target}}<0$. The target values for other constraints are set as follows: $\omega_{Y,\text{target}}=600~\mathrm{kHz}$, and $A_{\max}=50\%$. Based on these parameter settings, we compare the optimization results obtained from formulations (20) and (22) for $\gamma_{\mathrm{target}}=\pm 1\times 10^{-3}$.

### v2

@@ 2.3.2 Control of SN bifurcations on FRC
As discussed in Li et al. (2022); Liang et al. (2025), the SN bifurcations of the fixed points of the reduced dynamics (16) correspond to those of the periodic orbits of the full system (1). This correspondence allows us to control the occurrence of SN bifurcations in reduced coordinates. In this subsection, we derive the expression of the SN bifurcation points in the reduced coordinates and eliminate their occurrence by controlling the frequency separation between the two SN bifurcation points, namely, the $d_{\mathrm{SN}}$ shown in Fig. 1. In the final step, we formulate an optimization problem to suppress the occurrence of SN bifurcations.
We first need to find the fixed point of SN bifurcation $(\rho_{\mathrm{SN}},\Omega_{\mathrm{SN}})$ in reduced coordinates. The Jacobian of the reduced dynamics (16) can be expressed as Li (2024)
@@
At the SN bifurcation of a fixed point, the Jacobian $\boldsymbol{A}$ has a simple zero eigenvalue Kuznetsov (2023). Thus, the determinant of $\boldsymbol{A}$ is 0 at SN bifurcation, namely
@@
Note that there are usually two SN bifurcation points on the FRC, each corresponding to an excitation frequency at which an SN bifurcation occurs in the forced response of the system. By controlling the frequency separation between these two SN points, the SN bifurcation behavior can be effectively tailored, thereby regulating the associated jump phenomena in frequency sweeps. Next, we aim to suppress the occurrence of these two SN bifurcation points by controlling the magnitude of $d_{\text{SN}}:=|\Omega_{\mathrm{SN1}}-\Omega_{\mathrm{SN2}}|$ .
Many MEMS sensors (e.g., accelerometers, gyroscopes) are designed to operate within the linear regime to avoid hysteresis and excess noise caused by nonlinear effects Tiwari and Candler (2019). To ensure such behavior, we aim to control the system’s bifurcation structure while enhancing its performance. Specifically, we formulate an optimization problem that maximizes the response amplitude $\rho_{\max}$, subject to a constraint on the parameter $d_{\text{SN}}$ to limit the distance between the two SN bifurcation points. The resulting optimization problem is given as
@@
$\quad d_{\text{SN}}=|\Omega_{\text{SN1}}-\Omega_{\text{SN2}}|\leq d_{\text{target}}$
$\quad(\omega_{Y}/\omega_{Y,\text{target}}-1)^{2}\leq\varepsilon^{2}$
@@
where the constant $d_{\text{target}}$ regulates the distance between the two SN bifurcation points. As $d_{\text{target}}$ approaches zero, these two points coalesce, and the jump phenomenon is suppressed. The roles of the constraints on $\omega_{Y}$ and $\omega_{X}$ are discussed in Remark 4. A two-sided constraint is also imposed on the structural area to improve the stability of the optimization process.
Remark 4
We have consistently added the constraint $(\omega_{Y}/\omega_{Y,\text{target}}-1)^{2}\leq\varepsilon^{2}$ in the topology optimization problem formulations. This constraint enforces sufficient stiffness in the $Y$ direction throughout the optimization process, which is essential for obtaining a well-connected optimal topology layout. Indeed, we observe that the optimization iteration in (23) and (29) has difficulty converging to a well-defined optimal layout if this constraint is removed. We further add a constraint on $\omega_{X}$ to prevent the internal resonance between the master mode and other modes.
@@ 3 Sensitivity analysis
In the previous section, we formulated several optimization problems. Since the Method of Moving Asymptotes (MMA) is used to solve it, sensitivity analysis with respect to the design variables is essential for updating the design. In this section, we derive the sensitivity expressions for the key quantities $\omega_{j}$, $\lambda$, $\gamma$, $\tilde{f}$, $\rho_{\max}$, and $\Omega_{\text{SN}}$, which are critical for the optimization process.
The sensitivity of the eigenfrequencies $\omega_{j}$ and eigenvalues $\lambda$ has already been derived in Appendix.C of Li (2024), namely
@@
Following Li (2024), the variation of $\gamma$ can be obtained by the adjoint method, which can be expressed as
@@
In the above equation, the damping coefficients $\alpha$ and $\beta$ are fixed, while the damping ratio $\xi$ varies throughout the optimization process. This differs from the formulation in Li (2024), where the damping ratio remains constant. We follow the tensorial approach in Pozzi et al. (2025b) to ensure efficient implementation.
@@
Similarly, by taking derivative of (18), the sensitivity of $\Omega_{\text{SN}}$ can be expressed as
$\Omega_{\text{SN}}^{\prime}=c_{3}-\frac{c_{4}}{c_{5}}$ (34)
where
$c_{3}=$ $\mathrm{Im}(\lambda)^{\prime}+\mathrm{Im}(\gamma)^{\prime}\rho_{\text{SN}}^{2}+2\mathrm{Im}(\gamma)\rho_{\text{SN}}\rho_{\text{SN}}^{\prime},$ (35)
$c_{4}=$ $2\epsilon^{2}|\tilde{f}||\tilde{f}|^{\prime}-2(\mathrm{Re}(\lambda)\rho_{\text{SN}}+\mathrm{Re}(\gamma)\rho_{\text{SN}}^{3})$
$(\mathrm{Re}(\lambda)^{\prime}\rho_{\text{SN}}+\mathrm{Re}(\lambda)\rho_{\text{SN}}^{\prime}+\mathrm{Re}(\gamma)^{\prime}\rho_{\text{SN}}^{3}+3\mathrm{Re}(\gamma)\rho_{\text{SN}}^{2}\rho_{\text{SN}}^{\prime})$
$-2(\mathrm{Im}(\lambda)-\Omega_{\mathrm{SN}}+\mathrm{Im}(\gamma)\rho_{\text{SN}}^{2})^{2}\rho_{\text{SN}}\rho_{\text{SN}}^{\prime},$
$c_{5}=$ $2(\mathrm{Im}(\lambda)-\Omega_{\mathrm{SN}}+\mathrm{Im}(\gamma)\rho_{\text{SN}}^{2})\rho_{\text{SN}}^{2},$
and $\rho_{\mathrm{SN}}^{\prime}$ is given by Eq. (47) in Appendix B. One can further easily compute $d_{\text{SN}}^{\prime}$ via the sensitivity of $\Omega_{\text{SN}}^{\prime}$.
We conclude this section with an intuitive interpretation of how the sensitivities of $\rho_{\max}$, $\gamma$, and $d_{\text{SN}}$ guide the design evolution. As seen in Fig. 1, the parameter $\rho_{\max}$ corresponds to the peak amplitude of the FRC, and reducing $\rho_{\max}$ is generally associated with an increase in the effective stiffness of the structure. The parameter $\gamma$ characterizes the hardening or softening behavior of the nonlinear response. In particular, increasing $\mathrm{Im}(\gamma)$ weakens softening effects or enhances hardening behavior. Finally, the parameter $d_{\text{SN}}$ is related to the frequency range with the coexistence of stable and unstable periodic orbits on the FRC. A larger magnitude of $d_{\text{SN}}$ corresponds to a larger interval of unstable periodic orbits, leading to a more pronounced jump phenomenon in forward frequency sweep.
@@ 4 SIMP Model for Topology Optimization
In this work, the Solid Isotropic Material with Penalization (SIMP) method Bendsøe and Sigmund (1999) is employed as the foundation for density-based topology optimization. In the SIMP-based framework, the material properties of the structure are interpolated element-wise based on the physical density variables $\hat{\mu}_{e}$. These densities are used to assemble the global matrices and nonlinear force vectors. Specifically, the global mass and stiffness matrices, as well as the quadratic and cubic stiffness tensors, are computed as follows Pozzi et al. (2025b):
@@
To ensure mesh-independent and well-posed topology optimization results, the design variable $\mu\in[0,1]$ is first smoothed using a density filter Wang et al. (2011), defined as
@@
Next, a Heaviside-type projection Wang et al. (2011) is applied to obtain the projected density $\bar{\mu}_{e}$, which sharpens the transition between solid and void regions:
@@
Remark 5
The material distribution is initialized according to initial layouts (cf. Figs. 2, 6, and 9). Specifically, the design variables of the elements in the non-design region (black area) are fixed to $\mu_{e}=1$, whereas the design variables of the elements in the designable region (gray area) are uniformly initialized to the prescribed area fraction upper-bound $A_{\max}$, i.e., $\mu_{e}=A_{\max}$, such that the area fraction constraint is active.
@@ 5 Numerical examples
In this section, three numerical examples are presented to show the result of minimizing the peak of FRC, controlling hardening/softening behavior, and suppressing the occurrence of SN bifurcations. In all the examples, the 2-dimensional structures are discretized using 4-node, square elements. Following Pozzi et al. (2025b), material properties are specified with the density $\rho=2330\times 10^{-6}\ \rm{ng/um^{3}}$, Young’s modulus $E=148\times 10^{9}\ \rm{Pa}$, and Poisson’s ratio $\nu=0.23$. The plane stress approximation is used, with a thickness of 24 $\rm{\mu m}$.
The matrix assembly of the finite element model and the solution of the optimization problems are carried out using the YetAnotherFEcode package Jain et al. (2022). The FRCs of the optimized layouts are computed using SSMTool Jain et al. (2023), an open-source software package for the computation of invariant manifolds and their reduced dynamics. In the Sec. 2.2 and 2.3.1, we employed a third-order reduced-order model (16) to obtain the expression of FRC in the reduced coordinates (18). To assess the reliability of the optimized result, we further compute the FRC using SSMTool at higher-order truncations for the optimized structure. If the FRC at higher-order truncations agrees well with that at the third-order truncation, it indicates that the predicted FRC has converged and is hence reliable.
@@
We consider a Messerschmitt-Bölkow-Blohm (MBB) beam with a fixed segment located at its center. Due to structural symmetry, the optimization is performed only on the left half of the beam Pozzi et al. (2025b), as illustrated in Fig. 2. The half MBB beam measures 800 $\mathrm{\mu m}$ in length and 100 $\mathrm{\mu m}$ in height, including a fixed region with a length of 100 $\mathrm{\mu m}$. The left end of the beam is fully fixed, and the right end is constrained in the $x$-direction. A periodic excitation is applied at the midpoint of the right edge.
@@
With $\gamma_{\mathrm{target}}=5\times 10^{-5}$, the iterative process for the nonlinear optimization formulation (21) is provided in Fig. 3. The objective $\rho_{\max}$ exhibits an initial increase, followed by a rapid decrease until convergence. Meanwhile, the constraint on $\mathrm{Im}(\gamma)$ is violated initially, and the value of $\mathrm{Im}(\gamma)$ gradually converges and ultimately reaches the target $\mathrm{Im}(\gamma)=\gamma_{\mathrm{target}}$. The abrupt changes observed in the objective $\rho_{\max}$ and $\mathrm{Im}(\gamma)$ during the iteration are caused by the scheduled updates of the penalization exponent $p$ and the projection parameter $\sigma$ in the SIMP algorithm. Specifically, $\sigma$ is initialized at $\sigma_{0}=10$ and multiplied by $1.2$ every 40 iterations until it reaches a maximum value of $\sigma_{\max}=50$. Meanwhile, the parameter $p$ is initialized at $p_{0}=1$ and increased by one every 40 iterations, with an upper bound of $p_{\max}=8$. The obtained optimal layout is given in the upper panel of Fig. 4. The computational performance of this iterative process is summarized in Table 1Here and throughout this study, the stopping criterion for the optimization of $c_{\text{nl}}$ is defined by the simultaneous satisfaction of all constraints and the convergence of the objective function, i.e., $\max_{j\in\{i-3,i-2,i-1\}}\ |{c_{\text{nl}}^{(i)}-c_{\text{nl}}^{(j)}}|/{c_{\text{nl}}^{(i)}}<10^{-3}$. We check the convergence by a comparison with several previous iterations to ensure a steady convergence. In addition, the efficiency listed in all tables of this article is defined as the total optimization time divided by the number of optimization iterations, i.e., the average computational time per iteration, which includes the time to compute the SSM reduction, the objective function and also its sensitivity.
@@
Since third-order SSM reduction is employed to compute the FRCs of the optimal layouts, it is necessary to assess the convergence of the SSM-based model reduction. The convergence analysis is presented in Appendix C.1, which shows that the FRC obtained using third-order SSM reduction agrees well with that obtained using seventh-order SSM reduction under the forcing amplitude $f^{\mathrm{ext}}=2\times 10^{10}\ \mathrm{ng\cdot\mu m/ms^{2}}$.
To assess the robustness of the optimized structures, we vary both the forcing amplitude and the material stiffness and examine their effects on the resulting optimal layouts and dynamic responses. The corresponding optimization results are presented in Appendix C.2, which show that, within the range where the third-order SSM remains convergent, the optimized layouts and their dynamic responses show good robustness to these variations.
Our optimization framework assumes no internal resonance. To justify this assumption, the evolution of the first three eigenfrequencies during the iterative process for the case of $\gamma_{\mathrm{target}}=5\times 10^{-5}$, is presented in Appendix C.3. The results show that the natural frequencies are well separated and neither modal collisions nor internal resonances occur. As a result, we can track the target master mode in optimization easily by choosing the mode with lowest natural frequency. As detailed in Appendix C.3, the initial and optimized structures exhibit similar linear modal shapes for the master mode, which further confirms the consistency of master mode during optimization process.
The optimization results presented in this subsection show that the nonlinear optimization formulation in Eq. (21) achieves a minimization of the FRC peak comparable to that obtained with the linear optimization formulation in Eq. (22). This suggests that the SSM-based framework provides an effective approach for reducing the peak response of the FRC. More importantly, in contrast to the linear optimization formulation, the nonlinear framework allows the hardening or softening characteristics of the system to be actively tailored, highlighting an additional advantage of the SSM-based approach beyond peak-response reduction.
@@
Tailoring the hardening/softening behavior of the backbone curve in an MBB beam has been discussed in Pozzi et al. (2025b). In this subsection, we aim to minimize the response amplitude while simultaneously tailoring the softening–hardening behavior, namely, the concurrent design (21). We further compare this concurrent design with the design given by the reference problem (23) that tunes the backbone only. As shown in Fig. 6, the design domain of the beam has a length of 500 $\mathrm{\mu m}$ and a height of 100 $\mathrm{\mu m}$. A fixed region, occupying 20% of the total beam length, is located at the center and serves as a proof mass. The left end of the beam is fully fixed, while the right end is constrained in the $x$-direction. A periodic excitation is applied at the midpoint of the right edge.
@@
The FE model uses a mesh of $100\times 20$ elements, leading to a system with nn= 4179 DOF in (1). The damping ratio of initial layout $\xi$ = 0.1%, which leads to the values of damping constants $\alpha=6.90$ and $\beta=7.41\times 10^{-8}$. During the optimization iteration process, the forcing amplitude is given as $f^{\mathrm{ext}}=5\times 10^{9}\ \mathrm{ng\cdot\mu m/ms^{2}}$. In the following computation, we set $\epsilon=0.01$ unless otherwise stated. The projection parameter is fixed at $\sigma=10$ throughout the optimization, and the penalization parameter $p$ is initialized at $p_{0}=1$ and increased by one every 40 iterations, up to a maximum value of $p_{\max}=8$. The filter radius is set to $R=3$. In this example, no constraint is imposed on $\omega_{X}$ since $\omega_{X}>3\omega_{Y}$ is naturally satisfied throughout the optimization iterations. Furthermore, the two-sided constraint on $\gamma$ is replaced by a one-sided constraint. Specifically, $\gamma>\gamma_{\mathrm{target}}$ is enforced when $\gamma_{\mathrm{target}}>0$, and $\gamma<\gamma_{\mathrm{target}}$ is enforced when $\gamma_{\mathrm{target}}<0$. The target values for other constraints are set as follows: $\omega_{Y,\text{target}}=600~\mathrm{kHz}$, and $A_{\max}=50\%$. Based on these parameter settings, we compare the optimization results obtained from formulations (21) and (23) for $\gamma_{\mathrm{target}}=\pm 1\times 10^{-3}$.